CONFORMATIONS OF CATIONIZED PEPTIDES. DETERMINATION OF LIGAND BINDING GEOME-TRIES BY IRMPD SPECTROSCOPY

\underline{ROBERT C. DUNBAR}, Chemistry Department, Case Western Reserve Univ., Cleveland, OH 44106; JEFFREY STEILL, JOS OOMENS, FOM Institute for Plasma Physics, Nieuwegein, Netherlands; NICK C. POLFER, Chemistry Department, University of Florida, Gainesville, FL 32611.

Spectroscopic study of the conformations of metalated amino acids has mapped out in some detail the preferences for canonical (charge solvated) versus zwitterionic (salt bridge) conformations. Corresponding studies of larger peptides are now possible. Here are described results for several singly and doubly charged metal ions with dipeptides and tripeptides. Factors including ion charge, size of cation, and side chain identity and sequence are found to be conformational determinants. IRMPD spectra of the ions were acquired by irradiating the cell with infrared light from the FELIX free electron laser at wavelengths in the approximate range 500 to 1900 cm^{−1}.